# Processing-in-memory for genomics workloads

### v1

@@ Abstract
Figure 1: Typical workflow in genomics applications. DNA is extracted and then read using a sequencing instrument, which produces electrical resistance profiles, images, or “movies”. This signal data is converted into short DNA strings called reads in the basecalling step. Reads are then either aligned to an existing reference (resequencing), analyzed independently (alignment-free), or assembled into longer contiguous sequences (de novo sequencing).
@@
Analysis of genomic data involves many steps, most of which are computationally intensive. Figure 1 outlines a sequencing-based genomics workflow [2]. DNA is extracted, and a sequencing library is prepared. The instrument generates raw data (e.g., images, movies, or electrical resistance profiles), which base calling converts into DNA reads with quality scores. For simplicity, we categorize the sequence analysis into three classes: 1) resequencing when comparing to existing references, 2) alignment-free analysis, and 3) de novo assembly. Further analyses include variant effect prediction, genome annotation, and phylogenetics. Similar methods apply to RNA sequencing (transcriptomics[3].
Due to the large size of genomic data and the high computational requirements of algorithms in analyzing such data, several previous studies aimed to accelerate bioinformatics workloads using SIMD, FPGAs, and GPUs [4]. However, these solutions are still bottlenecked by data movement overhead. Processing-in-memory (PIM) technologies have emerged as promising approaches that fundamentally address the data movement bottleneck and provide effective parallelization capabilities. Transferring all data to a centralized processor - often located at a considerable distance from the storage system - leads to latency issues, bandwidth limitations, and high energy demands associated with data transfer and processing. On the other hand, PIM technologies facilitate computation directly at the location of the data. By eliminating the need for extensive data movement, PIM potentially achieves performance and energy efficiency improvements by several orders of magnitude when processing large-scale datasets.
We started the BioPIM Project (https://www.biopim.eu) to address computational challenges of genomics workloads, leveraging the advantages of PIM computing. BioPIM aims to realize cheap, ultra-fast, ultra-low energy processing that eliminates dependence on large, power-hungry computing clusters/data centers. One of the main goals of BioPIM is to characterize algorithms and data structures commonly used in bioinformatics, specifically algorithms for alignment and mapping, data structures such as Bloom Filters, hash tables, and graphs, as well as machine learning and deep learning methods. We have found that most data- and compute-intensive applications are memory bound, therefore good candidates for PIM acceleration (10.5281/zenodo.14858721). BioPIM’s research focuses on creating alternative PIM architecture designs for these applications and exploring their potential use in accelerating bioinformatics research, focusing on genomic sequence analysis.
@@
Emerging PnM solutions, such as SK Hynix’s AiM and Samsung’s HBM-PIM, embed processing units close to memory banks to improve performance and energy efficiency. Among these, UPMEM stands out as the only commercially available product at the time of preparing this manuscript. UPMEM integrates lightweight data processing units (DPUs) directly onto the DRAM die without any modifications to the manufacturing process. A typical configuration uses standard DDR4 DIMMs with 16 PIM chips. This architecture delivers thousands of DPUs at the server level, enabling local data processing, consequently dramatically reducing energy costs and data transfer overhead [14]. UPMEM’s robust software stack further simplifies application porting, making it a compelling solution for energy-sensitive, data-intensive workloads such as genomic analysis.
On the other hand, PuM uses analog properties of memory arrays (such as Kirchhoff laws or charge sharing) to implement computation within memory arrays directly on the bitlines, reading out the result via sense amplifiers or analog-to-digital converters (ADCs). This approach further alleviates the memory wall by considerably reducing or sometimes eliminating the need to transfer data outside memory arrays for processing. It also minimizes energy consumption by confining data movements within memory arrays. Moreover, PuM unlocks a very high degree of parallelism by enabling simultaneous processing, for example, in all memory columns in parallel. However, PuM is restricted due to the predefined topology and internal connectivity of memory arrays and can be efficiently used in embarrassingly parallel (e.g., element-wise or highly regular) processing tasks. Key computational workloads, including linear algebra operations and specific genome analyses like k-mer-based classification, exhibit properties well-suited for PuM. These tasks typically feature highly regular, embarrassingly parallel, SIMD-like computations on very large datasets, characteristics that map efficiently onto the inherent parallelism and structural regularity of memory-centric PuM systems.
@@
In this section, we present a high-level overview of PnM architectures developed within BioPIM for genomics applications. Our PnM designs implement similarity identification applications, both alignment-based and alignment-free, and variant calling. They target real-world PnM systems, i.e., they have been deployed on commercially available UPMEM platforms.
@@
Assessment of sequence similarity is a fundamental task in almost all bioinformatics pipelines. Similarities are identified through alignment algorithms, which are designed using dynamic programming (DP) and, therefore, have quadratic time and space complexity. In addition to computational challenges, sequence alignment remains a fundamentally memory-bounded computation with a low data reuse ratio. As a result, it suffers from constant data movement latency and a low memory bandwidth bottleneck. PnM architectures are strong candidates for accelerating alignment, thanks to their capacity to reduce data movement overhead and effectively parallelize numerous separate alignment tasks. In the following paragraphs, we present a high-level overview of deploying different alignment algorithms on UPMEM.
KSW2 library. Using the UPMEM PnM architecture, we have parallelized KSW2, a version of the Needleman-Wunsch algorithm (https://github.com/lh3/ksw2). KSW2 is intensively used in Minimap2, one of the most efficient software programs for DNA sequence alignment. We distribute sequence batches to each memory unit and perform alignment calculations independently using the computing resources attached to each memory unit. We evaluated our design using 9,500 16S RNA sequences and compared it with Minimap2. Our design demonstrates 9x speed-up on an UPMEM server with 160GB PnM memory while achieving $\sim$3.7x reduction in energy consumption compared to a server without PnM memory.
Smith-Waterman-Gotoh (SWG). We target the SWG algorithm, a DP-based local alignment method that incorporates an affine gap penalty model. SWG aims to identify similar regions between two sequences without constraints on their location. In our implementation, we evenly distribute independent sequence pairs across banks of different DPUs available in the UPMEM platform. A thread is launched in each DPU, and each DPU works independently to align a set of sequence pairs. We tested our implementation on an UPMEM system with 2560 DPUs (20 UPMEM-DIMMs) running at 425MHz and compared it with an OpenMP-optimized CPU baseline on server-grade CPU systems. Our PnM implementation delivered a 6x speedup over the CPU baseline running on an Intel Xeon Gold 5120 server for long-read datasets that strain the available memory resources.
@@
CGK embeddings [15] is a technique for approximating edit distance computations using Hamming distances, which can be calculated in linear time. Given the randomized nature of the algorithm, averaging multiple independent trials is necessary to reduce errors. We implemented CGK on UPMEM PnM architecture, and it is consistently 3x faster than its CPU counterpart for non-trivial situations where more than two sequences must be compared.
@@
Most read mappers use the seed-filter-extend strategy, where initial candidate positions are given by short, exact matches between the target sequence and the reference, which are then extended and optionally filtered in an iterative process to find the best overall alignment. The filtering step comes at the cost of reduced sensitivity since alternative alignments can be filtered out to favor the most likely solution.
A competing approach that tries to preserve all possibilities is the seed-embed-extend in which likely similar regions highlighted by seed matching are embedded into sketched representations. Edit distance estimations by comparing sketches are faster to compute than exact algorithms. All possible matching regions can then be ranked and reported to the user for a minimal loss of sensitivity.
Compression. Read mapping can also help improve data compression efficiency. Using PnM, we implemented a reference-guided sequence data compression algorithm. A typical sequencing experiment generates billions of short DNA reads, which are saved in files that are 100s of gigabytes. Rather than storing the reads, our algorithm stores their map locations and alignment differences relative to a reference genome. Therefore, the mapping stage is essential, accounting for most of the computational workload in the compression process. With only 4 UPMEM PnM modules, we achieved 2x speedup with 58% less energy over the state-of-the-art DNA read compression tool, Genozip, when run on a server with 16 CPU cores.
@@
Similarity searches can also be carried out without the need to calculate alignments. This is particularly needed when many genomes need to be analyzed, compared, or classified. These methods are based on the manipulation of k-mers, which are short sequences of $k$ nucleotides chosen to represent a genome. One of the preliminary steps in selecting k-mers is counting them, which typically requires sorting. The enumeration of k-mers in extensive sequencing datasets, as encountered in metagenomics (involving hundreds of billions of k-mers), presents a significant computational challenge. Consequently, substantial research efforts continue to be directed towards the optimization of computational time and memory utilization.
We implemented various sorting algorithms on the UPMEM PnM architecture to estimate the potential gain on the general k-mer counting problem. Despite our improvements, the resulting gains are comparatively small compared to SIMD acceleration on CPUs using the AVX instruction set.
@@
Bloom filters are space-saving probabilistic data structures for set membership queries. This data structure is commonly used for testing the existence of k-mers in collections of genomic sequences. Briefly, k-mers are “inserted” into a bit-vector using several hash functions, and then any k-mer can be queried in the structure. Bloom filters are also used in de Bruijn graph implementations to boost query efficiency. To improve the Bloom filter efficiency with potential effect on many applications, we have implemented it on the UPMEM PIM architecture. Our evaluation showed $\sim$3x speed up for inserting and querying 100 million elements in a Bloom filter of size 1 GB, compared to a traditional C++ implementation.
Perfect hash functions (PHFs) represent a specialized category of hash functions designed to eliminate collisions, ensuring that each unique key is mapped to a distinct index within the underlying data structure. This property enables constant-time ($O(1)$) lookup operations, significantly enhancing the computational efficiency of applications that rely on fast key retrieval. A minimal perfect hash function (MPHF) further constrains the output range such that the indices exactly span the size of the input key set. The minimum space required for such structures is known to be 1.44 bits per key. This efficiency stems from the fact that MPHFs assume a fixed, pre-defined set of keys, eliminating the need to store the keys themselves within the data structure — a key distinction from conventional hash tables.
@@
Among these, SicHash distinguishes itself through its highly parallelizable key placement strategy during construction. The algorithm partitions keys into uniformly sized buckets and subsequently employs cuckoo filters to establish the mapping between keys and their respective positions.
@@
Variant calling is a critical step in genome analysis, where mutations are identified by comparing the mapping information of reads generated from a sequenced genome to a reference. One of the most computationally intensive parts of this process is the Pair-Hidden Markov Model (Pair-HMM) algorithm, which calculates the probability of each read originating from each potential haplotype. This step has traditionally been a bottleneck on CPU or FPGA platforms due to extensive data movement and heavy floating-point calculations. Using the UPMEM PnM framework, these challenges can be overcome by moving the computation directly into memory, significantly reducing data transfer overhead and power consumption. In our approach (GAPiM) [8], we modified the Pair-HMM algorithm to operate in the logarithmic domain and use fixed-point arithmetic to replace slower multiplication and emulated floating-point operations. This innovation not only maintains acceptable levels of accuracy but also delivers up to 2x speedup compared to CPU SIMD implementations, and up to 3x speedup compared to FPGA solutions. These results demonstrate that even complex, data-intensive tasks such as variant calling can be efficiently performed on PIM platforms, ushering in a new era of scalable and energy-efficient genomic analysis.
@@
As illustrated in Figure 1, most bioinformatics pipelines begin by sequencing a genomic sample. This involves passing the sample through a flow cell sensor array and analyzing the DNA strands to infer the nucleotide sequence of each strand. While various techniques may be applied to perform this analysis, Oxford Nanopore Technology’s ultra-long read technology enables sequencing of DNA strands millions of bases long, reducing assembly complexity and improving accuracy on repetitive reads.
@@
Using DNNs for basecalling presents an opportunity for acceleration via PuM using non-volatile memory-based crossbar arrays. We performed extensive analysis and HW/SW co-optimization of ONT’s current basecaller Dorado (also known as Bonito) network for implementation on Phase Change Memory (PCM)-based PuM crossbars. Figure 3-b illustrates a complete PuM architecture for performing on-device basecalling entitled CiMBA (Compute-in-Memory Basecalling Accelerator) [16], measuring only 25mm2 and capable of being co-located at the data collection point.
The AL-Dorado network, illustrated in Figure 3-a, is, in turn, optimized for implementation on the PuM architecture by adjusting its layer sizes to maximally utilize the available PuM resources and is retrained to be robust against PCM nonidealities and noise sources using the popular AIHWKIT library [17]. We validate our network retraining accuracy on real PCM-based crossbar arrays, which allows us to quantitatively evaluate the impact of PuM non-idealities. We find that implementing the network in PCM arrays without retraining results in a 7.66% loss in accuracy, while retaining the network for resistance against noise reduces this accuracy drop to only 1.96%.
We tested the entire system in a PuM-based simulator to analyze runtime latency, energy, and area performance of DNNs on PuM crossbar architectures. We find that, due to the massive parallelism and pipelining capabilities provided by PuM-based architectures, we can achieve a throughput 2x faster than SotA-embedded GPU basecalling solutions and 24x the necessary throughput to achieve real-time basecalling for portable ONT sequencing devices. The architecture is also 16.5x more power-efficient against SotA embedded GPUs, improving field utility by enabling energy-limited setups such as those powered by batteries to operate for longer periods.
@@
Accurate classification of pathogen genomes from unstructured and sparse sequenced data is essential for disease surveillance, outbreak tracking, and antimicrobial resistance monitoring. Pathogen classification identifies virulent strains, pathogen transmission dynamics, and genetic variations contributing to pathogenicity. With the increasing use of metagenomics, large-scale sequencing datasets require robust computational methods to classify pathogens efficiently.
Our classification platform applies k-mer matching, a heuristic that queries newly sequenced DNA fragments in a reference database. Our unique solution implements pathogen genome classification in a special class of memory - a similarity search capable Content Addressable Memory (SAS-CAM).
GCOC [10], developed within BioPIM, is a fast and highly sensitive approximate k-mer matching-based on-chip genome detection and classification platform. GCOC SoC architecture, data flow, and evaluation setup are presented in Figure 4. GCOC comprises SAS-CAM, a dual-issue RISC-V CPU, and a Streaming Buffer (SBUFF), which allows quick uploading to and updating of the reference database in SAS-CAM. GCOC SoC has been manufactured using a commercial $65nm$ process.
@@
SAS-CAM is a 9T NOR CAM with a matchline shared across each row and row sense amplifiers. Its key distinction is an added NMOS transistor in each CAM cell, controlling matchline discharge speed. The similarity tolerance threshold is set by adjusting this transistor’s gate voltage and the row sense amplifiers’ reference voltage, enabling dynamic control over classifier sensitivity, specificity, and precision.
Figure 4 illustrates genome classification in GCOC. The reference genome database (k-mers of predefined organisms) is uploaded offline to SAS-CAM through SBUFF, which RISC-V feeds slower. During classification, RISC-V fetches newly sequenced DNA reads and extracts from them k-mers as queries in SAS-CAM. The pointers to the matching k-mers are fed back to RISC-V, which associates them with known taxa.
@@
GCOC evaluation setup, test board, SoC, and SAS-CAM layout are shown in Figure 4. GCOC processing-using-CAM design enables low power operation (only 1.27mW during similarity search) and throughput of up to 400 M k-mers/sec at 100MHz. Our on-silicon experiments show that GCOC can reach a sensitivity of 96% and specificity of 100% in bacterial and viral genome classification with the Hamming distance threshold set to 3.
While GCOC uses SRAM-based CAM, alternative memory infrastructures exist. DASH-CAM [12] employs Gain Cell eDRAM for k-mer-based genome classification, expanding the on-chip reference database by 5.5x. DIPER leverages resistive RAM (ReRAM) [11] for greater capacity and supports both Hamming and edit distance tolerance, improving accuracy. Finally, Major-K [13] utilizes commodity DRAM for large-scale, cost-effective, and accessible k-mer matching.

### v2

@@ Abstract
Figure 1: Typical workflow in genomics applications. DNA is extracted and then sequenced using a sequencing instrument, which produces electrical resistance profiles, images, or “movies”. This signal data is converted into short DNA sequences, called reads, during the basecalling step. Reads are then either aligned to an existing reference (resequencing), analyzed independently (alignment-free), or assembled into longer contiguous sequences (de novo sequencing).
@@
Analysis of genomic data involves many steps, most of which are computationally intensive. Figure 1 outlines a sequencing-based genomics workflow [2]. DNA is extracted and prepared for sequencing. The instrument generates raw data (e.g., images, movies, or electrical resistance profiles), which base calling converts into DNA reads with quality scores. For simplicity, we categorize the sequence analysis into three classes: 1) resequencing when comparing to existing references, 2) alignment-free analysis, and 3) de novo assembly. Further analyses include variant effect prediction, genome annotation, and phylogenetics. Similar methods apply to RNA sequencing (transcriptomics[3].
Due to the large size of genomic data and the high computational requirements of algorithms in analyzing such data, several previous studies aimed to accelerate bioinformatics workloads using SIMD, FPGAs, and GPUs, such as SneakySnake [4], Dragen, and ParaBricks, as well as entire analysis pipelines with GPU acceleration (urWGA). However, these solutions are still bottlenecked by data movement overhead. Processing-in-memory (PIM) technologies have emerged as promising approaches that fundamentally address the data movement bottleneck and enable effective parallelization. Transferring all data to a centralized processor - often located at a considerable distance from the storage system - leads to latency issues, bandwidth limitations, and high energy demands associated with data transfer and processing. On the other hand, PIM technologies facilitate computation at the data’s location. By eliminating the need for extensive data movement, PIM can achieve performance and energy-efficiency improvements by several orders of magnitude when processing large-scale datasets.
We started the BioPIM Project (https://www.biopim.eu) to address computational challenges of genomics workloads, leveraging the advantages of PIM computing. BioPIM aims to realize cheap, ultra-fast, ultra-low energy processing that eliminates dependence on large, power-hungry computing clusters/data centers. One of the main goals of BioPIM is to characterize algorithms and data structures commonly used in bioinformatics, including alignment and mapping algorithms, data structures such as Bloom Filters, hash tables, and graphs, and machine learning and deep learning methods. We have found that most data- and compute-intensive applications are memory-bound, therefore good candidates for PIM acceleration (10.5281/zenodo.14858721). BioPIM’s research focuses on developing alternative PIM architecture designs for these applications and exploring their potential to accelerate genomic sequence analysis.
@@
Emerging PnM solutions include SK Hynix’s AiM, Samsung’s HBM-PIM, and UPMEM [14]. Each of these embed processing units close to memory banks to improve performance and energy efficiency. Among these, UPMEM stands out as the only commercially available product at the time of preparing this manuscript. UPMEM integrates lightweight data processing units (DPUs) directly onto the DRAM die without any modifications to the manufacturing process. A typical configuration uses standard DDR4 DIMMs with 16 PIM chips. This architecture delivers thousands of DPUs at the server level, enabling local data processing, consequently dramatically reducing energy costs and data transfer overhead [14]. UPMEM’s robust software stack further simplifies application porting, making it a compelling solution for energy-sensitive, data-intensive workloads such as genomic analysis. However, the current generation of UPMEM DPUs do not support floating-point operations natively; therefore, they require emulated floating-point operations.
On the other hand, PuM uses analog properties of memory arrays (such as Kirchhoff laws or charge sharing) to implement computation within memory arrays directly on the bitlines, reading out the result via sense amplifiers or analog-to-digital converters (ADCs). This approach further alleviates the memory constraints by considerably reducing or sometimes eliminating the need to transfer data outside memory arrays for processing. It also minimizes energy consumption by confining data movements within memory arrays. Moreover, PuM unlocks a very high degree of parallelism by enabling simultaneous processing, for example, in all memory columns in parallel. PnM places processing units within the memory chip but outside memory arrays. Hence, it allows for very high bandwidth (due to a wide access bus) and reduces energy consumption (due to the close proximity of processing units to memory). However, PuM improves energy efficiency significantly beyond that by confining computing to the memory arrays themselves, within the frame of the sense amplifiers. A 2-3 orders of magni- tude energy growth has been reported between the sense amplifiers’ edge and the CPU. PuM potentially eliminates this energy growth. Key computational workloads, including linear algebra operations and specific genome analyses like k-mer-based classification, exhibit properties well-suited for PuM. These tasks typically feature highly regular, embarrassingly parallel, SIMD-like computations on very large datasets, characteristics that map efficiently onto the inherent parallelism and structural regularity of memory-centric PuM systems.
@@
In this section, we present a high-level overview of PnM architectures developed within BioPIM for genomics applications. Our PnM designs implement applications for similarity identification, both alignment-based and alignment-free, and for variant calling. They target real-world PnM systems, i.e., they have been deployed on commercially available UPMEM platforms.
@@
Assessment of sequence similarity is a fundamental task in almost all bioinformatics pipelines. Similarities are identified using alignment algorithms, which are based on dynamic programming (DP) and therefore have quadratic time and space complexity. In addition to computational challenges, sequence alignment remains a fundamentally memory-bound computation with a low data reuse ratio. As a result, it suffers from constant data movement latency and a low memory bandwidth bottleneck. PnM architectures are strong candidates for accelerating alignment, thanks to their ability to reduce data movement overhead and effectively parallelize multiple alignment tasks. In the following paragraphs, we present a high-level overview of deploying different alignment algorithms on UPMEM.
KSW2 library. Using the UPMEM PnM architecture, we have parallelized KSW2, a version of the Needleman-Wunsch algorithm (https://github.com/lh3/ksw2). KSW2 is intensively used in Minimap2, one of the most efficient software programs for DNA sequence alignment. We distribute sequence batches to each memory unit and perform alignment calculations independently using the computing resources attached to it. We evaluated our design using 9,500 16S RNA sequences and compared it with Minimap2. Our design demonstrates 9x speed-up on an UPMEM server with 160GB PnM memory while achieving $\sim$3.7x reduction in energy consumption compared to a server without PnM memory.
Smith-Waterman-Gotoh (SWG). We target the SWG algorithm, a DP-based local alignment method that incorporates an affine gap penalty model. SWG aims to identify similar regions between two sequences without constraints on their location. In our implementation, we evenly distribute independent sequence pairs across banks of different DPUs available in the UPMEM platform. A thread is launched in each DPU, and each DPU works independently to align a set of sequence pairs. We tested our implementation on a UPMEM system with 2560 DPUs (20 UPMEM-DIMMs) running at 425 MHz and compared it with an OpenMP-optimized CPU baseline on server-grade CPU systems. For aligning long-read datasets that strain available memory, our PnM implementation delivers up to 6x speedup over a CPU baseline on an Intel Xeon Gold 5120 server.
@@
CGK embeddings [15] is a technique for approximating edit distance computations using Hamming distances, which can be calculated in linear time. Given the randomized nature of the algorithm, averaging multiple independent trials is necessary to reduce errors. We implemented CGK on the UPMEM PnM architecture, and it is consistently 3x faster than its CPU counterpart in non-trivial situations where more than 2 sequences must be compared.
@@
Most read mappers use the seed-filter-extend strategy, where initial candidate positions are determined by short, exact matches between the target sequence and the reference, which are then extended and optionally filtered iteratively to find the best overall alignment. The filtering step comes at the cost of reduced sensitivity since alternative alignments can be filtered out to favor the most likely solution.
A competing approach that tries to preserve all possibilities is the seed-embed-extend in which likely similar regions highlighted by seed matching are embedded into sketched representations. Edit distance estimations based on sketch comparisons are faster to compute than exact algorithms. All possible matching regions can then be ranked and reported to the user for a minimal loss of sensitivity.
Compression. Read mapping can also help improve data compression efficiency. Using PnM, we implemented a reference-guided algorithm for compressing sequence data. A typical sequencing experiment generates billions of short DNA reads, which are stored in files that are 100s of gigabytes in size. Rather than storing the reads, our algorithm stores their map locations and alignment differences relative to a reference genome. Therefore, the mapping stage is essential, accounting for most of the computational workload in the compression process. With only 4 UPMEM PnM modules, we achieved 2x speedup and 58% lower energy consumption than the state-of-the-art DNA read compression tool, Genozip, when run on a server with 16 CPU cores.
@@
Similarity searches can also be performed without calculating alignments. This is particularly needed when many genomes need to be analyzed, compared, or classified. These methods are based on the manipulation of k-mers, which are short sequences of $k$ nucleotides chosen to represent a genome. One of the preliminary steps in selecting k-mers is counting them, which typically requires sorting. The enumeration of k-mers in extensive sequencing datasets, as encountered in metagenomics (involving hundreds of billions of k-mers), presents a significant computational challenge. Consequently, substantial research efforts continue to be directed toward optimizing computational time and memory utilization.
We implemented various sorting algorithms on the UPMEM PnM architecture to estimate the potential gain on the general k-mer counting problem. Despite our improvements, the resulting gains are comparatively small compared to those from SIMD acceleration on CPUs with the AVX instruction set.
@@
Bloom filters are space-saving probabilistic data structures for set membership queries. This data structure is commonly used to test for the presence of k-mers in collections of genomic sequences. Briefly, k-mers are “inserted” into a bit-vector using several hash functions, and then any k-mer can be queried in the structure. Bloom filters are also used in de Bruijn graph implementations to boost query efficiency. To improve Bloom filter efficiency and its potential impact across many applications, we have implemented it on the UPMEM PIM architecture. Our evaluation showed $\sim$3x speed up for inserting and querying 100 million elements in a Bloom filter of size 1 GB, compared to a traditional C++ implementation.
Perfect hash functions (PHFs) are a specialized category of hash functions designed to eliminate collisions, ensuring that each unique key maps to a distinct index within the underlying data structure. This property enables constant-time ($O(1)$) lookup operations, significantly enhancing the computational efficiency of applications that rely on fast key retrieval. A minimal perfect hash function (MPHF) further constrains the output range such that the indices exactly span the size of the input key set. The minimum space required for such structures is known to be 1.44 bits per key. This efficiency stems from the fact that MPHFs assume a fixed, predefined set of keys, eliminating the need to store the keys themselves within the data structure—a key distinction from conventional hash tables.
@@
Among these, SicHash distinguishes itself by its highly parallelizable key-placement strategy during construction. The algorithm partitions keys into uniformly sized buckets and subsequently uses cuckoo filters to map each key to its position.
@@
Variant calling is a critical step in genome analysis, where mutations are identified by comparing the mapping information of reads generated from a sequenced genome to a reference. One of the most computationally intensive parts of this process is the Pair-Hidden Markov Model (Pair-HMM) algorithm, which calculates the probability of each read originating from each potential haplotype. This step has traditionally been a bottleneck on CPU- or FPGA-based platforms due to extensive data movement and heavy floating-point computation. Using the UPMEM PnM framework, these challenges can be overcome by moving the computation directly into memory, significantly reducing data transfer overhead and power consumption. In our approach (GAPiM) [8], we modified the Pair-HMM algorithm to operate in the logarithmic domain and use fixed-point arithmetic to replace slower multiplication and emulated floating-point operations. This innovation not only maintains acceptable levels of accuracy but also delivers up to 2x the speed of CPU SIMD implementations and up to 3x the speed of FPGA solutions. These results demonstrate that even complex, data-intensive tasks such as variant calling can be performed efficiently on PIM platforms, ushering in a new era of scalable, energy-efficient genomic analysis.
@@
As illustrated in Figure 1, most bioinformatics pipelines begin by sequencing a genomic sample. This involves passing the sample through a flow-cell sensor array and analyzing the DNA strands to infer the nucleotide sequence of each. While various techniques may be applied to perform this analysis, Oxford Nanopore Technology’s ultra-long read technology enables sequencing of DNA strands millions of bases long, reducing assembly complexity and improving accuracy on repetitive reads.
@@
Using DNNs for basecalling offers an opportunity to accelerate PuM via non-volatile memory-based crossbar arrays. We performed extensive analysis and HW/SW co-optimization of ONT’s current basecaller Dorado (also known as Bonito) network for implementation on Phase Change Memory (PCM)-based PuM crossbars. Figure 3-b illustrates a complete PuM architecture for performing on-device basecalling entitled CiMBA (Compute-in-Memory Basecalling Accelerator) [16], measuring only 25mm2 and capable of being co-located at the data collection point.
The AL-Dorado network, illustrated in Figure 3-a, is, in turn, optimized for implementation on the PuM architecture by adjusting its layer sizes to maximally utilize the available PuM resources and is retrained to be robust against PCM nonidealities and noise sources using the popular AIHWKIT library [17]. We validate our network retraining accuracy on real PCM-based crossbar arrays, which allows us to quantitatively evaluate the impact of PuM non-idealities. We find that implementing the network in PCM arrays without retraining results in a 7.66% loss in accuracy, whereas retaining the network to reduce noise reduces this drop to only 1.96%.
We tested the entire system in a PuM-based simulator to analyze runtime latency, energy, and area performance of DNNs on PuM crossbar architectures. We find that, due to the massive parallelism and pipelining capabilities provided by PuM-based architectures, we can achieve a throughput 2x that of SotA-embedded GPU-based basecalling solutions and 24x the required throughput for real-time basecalling on portable ONT sequencing devices. The architecture is also 16.5x more power-efficient than SotA embedded GPUs, improving field utility by enabling energy-limited setups, such as those powered by batteries, to operate for longer periods.
@@
Accurate classification of pathogen genomes from unstructured, sparse sequencing data is essential for disease surveillance, outbreak tracking, and antimicrobial resistance monitoring. Pathogen classification identifies virulent strains, pathogen transmission dynamics, and genetic variations contributing to pathogenicity. With the increasing use of metagenomics, large-scale sequencing datasets require robust computational methods to efficiently classify pathogens.
Our classification platform uses k-mer matching, a heuristic that queries a reference database for newly sequenced DNA fragments. Our unique solution implements pathogen genome classification in a special class of memory - a similarity search capable Content Addressable Memory (SAS-CAM).
@@
GCOC [10], developed within BioPIM, is a fast and highly sensitive approximate k-mer matching-based on-chip genome detection and classification platform. GCOC SoC architecture, data flow, and evaluation setup are presented in Figure 4. GCOC comprises SAS-CAM, a dual-issue RISC-V CPU, and a Streaming Buffer (SBUFF), which enables quick uploads and updates to the reference database in SAS-CAM. GCOC SoC has been manufactured using a commercial $65nm$ process.
SAS-CAM is a 9T NOR CAM with a matchline shared across each row and row sense amplifiers. Its key distinction is an additional NMOS transistor in each CAM cell, which controls the matchline discharge speed. The similarity tolerance threshold is set by adjusting this transistor’s gate voltage and the row sense amplifiers’ reference voltage, enabling dynamic control over classifier sensitivity, specificity, and precision.
Figure 4 illustrates genome classification in GCOC. The reference genome database (k-mers of predefined organisms) is uploaded offline to SAS-CAM through SBUFF, which RISC-V feeds more slowly. During classification, RISC-V fetches newly sequenced DNA reads and extracts k-mers from them as queries for SAS-CAM. The pointers to the matching k-mers are fed back to RISC-V, which associates them with known taxa.
@@
GCOC evaluation setup, test board, SoC, and SAS-CAM layout are shown in Figure 4. GCOC processing using CAM design enables low-power operation (only 1.27 mW during similarity search) and throughput of up to 400 M k-mers/sec at 100 MHz. Our on-silicon experiments show that GCOC can achieve 96% sensitivity and 100% specificity in bacterial and viral genome classification with a Hamming distance threshold of 3.
While GCOC uses SRAM-based CAM, alternative memory infrastructures exist. DASH-CAM [12] employs Gain Cell eDRAM for k-mer-based genome classification, expanding the on-chip reference database by 5.5x. DIPER leverages resistive RAM (ReRAM) [11] for greater capacity and support for both Hamming and edit distance tolerances, improving accuracy. Finally, Major-K [13] utilizes commodity DRAM for large-scale, cost-effective, and accessible k-mer matching.